OpenStudy (anonymous):

Help! Please and thank you. Solve for a 5(a - 4) - 8a = 55

OpenStudy (anonymous):

a = -25 (did you want a walkthrough?)

OpenStudy (anonymous):

Thanks, and sure!

OpenStudy (anonymous):

okay, gimme on second & no probelm

OpenStudy (anonymous):

And, I would love your help with more questions if you have the time.

OpenStudy (anonymous):

Okay 5(a - 4) - 8a = 55 then you distribute the 5 to the (a - 4) 5a - 20 - 8a = 55 then you add 5a to -8a -3a - 20 = 55 add the 20 to the 55 -3a = 75 then divide the -3 into 75 a = -25 & sure, I have time :)

OpenStudy (anonymous):

I'll even help out if you would like

OpenStudy (anonymous):

Awesome, Thanks! Here is the next one... The length of a rectangle is 12 in. more than its width. The perimeter is 64 in. What is the length of the rectangle?

OpenStudy (anonymous):

Gimme a second to figure it out :)

OpenStudy (anonymous):

Same

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

sosotiff i'll let you take over the bell just rang

OpenStudy (anonymous):

Thank anyway @Kira96 !

OpenStudy (anonymous):

okay ^-^ @Kira96

OpenStudy (anonymous):

You still there?

OpenStudy (anonymous):

Okay here's the answer; To find the perimeter you use 2(length+width). We'll use x to represent the width and x + 12 to represent the length. So, 2(x + x + 12) = 64 add inside the parentheses 2( 2x + 12) = 64 distribute the 2 4x + 24 = 64 subtract the 24 4x = 44 then divide the 4 x = 11 Then we plug it in 11 + 12 = 23 So the length is 23 inches

OpenStudy (anonymous):

Thank you very much, Im turning in my test now :)

OpenStudy (anonymous):

You're welcome, glad to help :)